OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ n! }{ (n+5)! }\] if i use the ratio test on this is the answer 1/5?

OpenStudy (anonymous):

the answer i get is 1

OpenStudy (anonymous):

\[\frac{ (n+1)! / (n+6)! }{ (n+5)!/n! } = \frac{ (n+1)*n! /n!*(n+1)(n+2)... (n+5)(n+6) }{ (n+1)(n+2)...(n+5)*n!/n! }\] taking the limit as n approaches infinity = 1

OpenStudy (anonymous):

some cut off

OpenStudy (anonymous):

\[\frac{ n+1 }{ n+6 }\]

OpenStudy (anonymous):

I trust your answer more than mine, it looks like you know your stuff better ;)

OpenStudy (anonymous):

hehe :) i love calculus

OpenStudy (anonymous):

:( calc kills me

OpenStudy (anonymous):

hehe. it means no harm :)